OpenStudy (anonymous):

solve using the quadratic formula x^2=4x+21

OpenStudy (anonymous):

Simplifying x2 = 4x + 21 Reorder the terms: x2 = 21 + 4x Solving x2 = 21 + 4x Solving for variable 'x'. Reorder the terms: -21 + -4x + x2 = 21 + 4x + -21 + -4x Reorder the terms: -21 + -4x + x2 = 21 + -21 + 4x + -4x Combine like terms: 21 + -21 = 0 -21 + -4x + x2 = 0 + 4x + -4x -21 + -4x + x2 = 4x + -4x Combine like terms: 4x + -4x = 0 -21 + -4x + x2 = 0 Factor a trinomial. (-3 + -1x)(7 + -1x) = 0 Subproblem 1Set the factor '(-3 + -1x)' equal to zero and attempt to solve: Simplifying -3 + -1x = 0 Solving -3 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '3' to each side of the equation. -3 + 3 + -1x = 0 + 3 Combine like terms: -3 + 3 = 0 0 + -1x = 0 + 3 -1x = 0 + 3 Combine like terms: 0 + 3 = 3 -1x = 3 Divide each side by '-1'. x = -3 Simplifying x = -3 Subproblem 2Set the factor '(7 + -1x)' equal to zero and attempt to solve: Simplifying 7 + -1x = 0 Solving 7 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-7' to each side of the equation. 7 + -7 + -1x = 0 + -7 Combine like terms: 7 + -7 = 0 0 + -1x = 0 + -7 -1x = 0 + -7 Combine like terms: 0 + -7 = -7 -1x = -7 Divide each side by '-1'. x = 7 Simplifying x = 7Solutionx = {-3, 7}

OpenStudy (cwrw238):

x =[ --4 +- sqrt(16 - 4x1x-21)] / 2 =( 4 + 10) / 2 = 7 also (4-10) / 2 = -3 solution set is {-3,7}

hero (hero):

Did somebody just copy and paste bagatrix?

OpenStudy (anonymous):

Actually its http://www.algebrahelp.com/

hero (hero):

Same thing